The common stock of Jensen Shipping has an expected return of 17.10 percent. The return on the market is 12 percent and the risk-free rate of return is 4.5. What is the beta of this stock?

options: 1.26 1.68 0.75 1.41 1.52

A stock has an expected return of 12 percent, the risk-free rate is 5.4 percent, and the market risk premium is 5 percent. The beta of this stock must be

options: 2.78 1.10 1.48 1.32 3.44
Answer #1

a.Expected return=risk free rate+beta*(market rate-risk free rate)

17.1=4.5+Beta*(12-4.5)

Beta=(17.1-4.5)/(12-4.5)

=1.68

b.Expected rate=risk free rate+Beta*market risk premium

12=5.4+(5*Beta)

Beta=(12-5.4)/5

=1.32